OpenStudy (anonymous):

A block of mass m1 = 5 kg rests on a table. Two ropes connect that block to two masses that hang off the sides of the table. The coefficient of static friction µs between the block and the table is 0.2. If one of the hanging masses has mass m2 = 3 kg, for what range of values of the other block’s mass m3 will the system not move?

OpenStudy (shamim):

Limiting static force F1=coefficient of static friction*m*g=0.2*5*9.8=? Weight of 3kg F2=mg=3*9.8=? Let wight of m3 mass is F3 Now the difference between F2 nd F3 will b equal to or less than F1

OpenStudy (shamim):

Feel free to ask for more clarification